# Power Battery Detection

### v1

@@ Abstract
Power batteries are essential components in electric vehicles, where internal structural defects can pose serious safety risks. We conduct a comprehensive study on a new task, power battery detection (PBD), which aims to localize the dense endpoints of cathode and anode plates from industrial X-ray images for quality inspection. Manual inspection is inefficient and error-prone, while traditional vision algorithms struggle with densely packed plates, low contrast, scale variation, and imaging artifacts. To address this issue and drive more attention into this meaningful task, we present PBD5K, the first large-scale benchmark for this task, consisting of 5,000 X-ray images from nine battery types with fine-grained annotations and eight types of real-world visual interference. To support scalable and consistent labeling, we develop an intelligent annotation pipeline that combines image filtering, model-assisted pre-labeling, cross-verification, and layered quality evaluation. We formulate PBD as a point-level segmentation problem and propose MDCNeXt, a model designed to extract and integrate multi-dimensional structure clues including point, line, and count information from the plate itself. To improve discrimination between plates and suppress visual interference, MDCNeXt incorporates two state space modules. The first is a prompt-filtered module that learns contrastive relationships guided by task-specific prompts. The second is a density-aware reordering module that refines segmentation in regions with high plate density. In addition, we propose a distance-adaptive mask generation strategy to provide robust supervision under varying spatial distributions of anode and cathode positions. Without any bells and whistles, our segmentation-based MDCNeXt consistently outperforms various other corner detection, crowd counting and general/tiny object detection-based solutions, making it a strong baseline that can help facilitate future research in PBD. Finally, we provide a detailed discussion of future directions and challenges for research in power battery detection. The source code and datasets will be publicly available at PBD5K.
Index Terms:
Power Battery Detection, Benchmark, Point-level Segmentation, Multi-dimensional Clues, State Space Modeling.
publicationid: pubid:
I Introduction
@@
With the rapid adoption of electric vehicles (EVs), Fig. LABEL:fig:development_EV(a) shows that EV market share has grown from around 1% to over 20% in the past five years, with 17.1 million units sold globally in 2024 [1, 2]. The power battery serves as the sole energy source for EVs, directly influencing their performance, range, and safety [3]. As shown in Fig. LABEL:fig:development_EV(b), the battery pack contributes nearly 35% of total manufacturing cost, underscoring its critical economic and functional role. Ensuring its structural integrity requires reliable inspection throughout production. During cell assembly (Fig. LABEL:fig:development_EV(c)), hundreds of cathode and anode plates are stacked alternately. Any misalignment or overhang can lead to short circuits, overheating, or even explosion, creating serious safety risks. To address this, a post-assembly inspection process, termed power battery detection (PBD), is widely applied. As shown in Fig. LABEL:fig:development_EV(d), PBD uses X-ray images captured by digital radiography (DR) systems to localize cathode and anode endpoints with coordinate-level precision. This enables automatic computation of key quality indicators, such as the number of plates and their overhang [^1] Number and overhang are two critical metrics in battery quality control. Number indicates the total count of anode/cathode plates, while overhang measures the average vertical offset between one anode and its adjacent cathodes., which determine whether a battery cell is acceptable (OK) or defective (NG). However, long-term manual inspections are subject to visual fatigue and human errors. Even top manufacturers relying on large re-inspection teams face high labor costs and limited efficiency.
@@
Figure 1: Visual examples from the PBD5K dataset (best viewed zoomed in). (a) 1-pixel localization on high-resolution (3000 × 4000) X-ray images. (b) The imaging field-of-view include Close Shot (CS), Medium Shot (MS) and Long Shot (LS). (c) Examples of various attributes (best viewed zoomed in). See Tab. I for details.
TABLE I: Attribute descriptions (see examples in Fig. 1(c)).
@@
Therefore, the development of an intelligent PBD model is an imminent need. To benchmark this new task, we first introduce PBD5K, the first large-scale power battery dataset with 5,000 X-ray images. It is constructed using an intelligent annotation pipeline that combines automatic filtering, model-assisted labeling, cross-validation, and hierarchical quality control, significantly improving labeling efficiency and accuracy. Compared with traditional visual tasks, PBD presents several unique challenges: I) Open Vision Problem Modeling. Like emerging tasks such as character stroke extraction [4], insubstantial object detection [5], open vocabulary objects detection and segmentation [6, 7], how to model a suitable AI-based solution is critical for the PBD task. II) Single Pixel-Level Object Localization. As shown in Fig. 1(a), PBD requires accurately identifying plate endpoints that are only one pixel in size among millions of pixels, and providing their precise coordinates. This level of search difficulty is akin to finding a needle in a haystack. III) Multi-shot Imaging. As shown in Fig. 1(b), X-ray imaging is conducted under various field-of-view settings, leading to significant variations in spatial resolution. Even under the same imaging setup, batteries may exhibit different physical dimensions or placement scales. IV) Weak Feature Perception. As clearly observed in Fig. 1(c), the plate endpoints exhibit extremely weak features and are visually similar to their surrounding regions. Especially under interference from bifurcated plates, separators, or trays, their visibility is further diminished. V) Fine-grained Structural Semantics. PBD5K contains a wide range of overhang distribution patterns. When the distance between adjacent cathode and anode plates is small (less than 100 pixels), their visual appearances become easily confused. Since power batteries are manufactured with strict alternation between cathode and anode plates, accurate endpoint localization implicitly requires the model to understand the structural semantics, including the classification and ordering of electrode plates.
@@
In this paper, we propose a multi-dimensional collaborative network for accurate PBD. Imitating the coarse-to-fine human visual perception system [8, 9], the model first perceives the battery region and then gradually refines the localization of each plate. Once the endpoints are precisely segmented, both their coordinates and total numbers can be derived. First, we reformulate PBD as a point-centric segmentation task, where the core objective is to predict a point map of single-pixel endpoints for cathode and anode plates. A U-shaped encoder-decoder [10, 11] serves as the basic structure for point prediction. Inspired by multi-task learning [12, 13, 14] and the structural traits of power batteries, we introduce two auxiliary tasks: line segmentation and plate counting. As shown in Fig. 2, connecting predicted endpoints yields line maps that constrain spatial position and reduce ambiguity. Line supervision guides endpoint refinement, while the counting branch enforces consistency between the predicted and actual number of electrodes, mitigating false positives and omissions. By jointly learning from these three complementary clues, we establish a segmentation-centered multi-dimensional collaborative framework. Next, to improve localization under weak textures and noisy backgrounds, we utilize state space models (SSMs) [15, 16, 17, 18], which offer efficient long-range modeling with linear complexity. Compared to Transformer-based models [19, 20], SSMs are more suitable for high-resolution industrial images due to their lower computational cost. However, direct use of SSMs may amplify irrelevant large-region patterns (e.g., trays or tabs), thereby overwhelming the subtle and sparse endpoint signals. To address this, we propose the prompt-filtered state space module (PFSSM). It uses features from a representative prompt image to generate dynamic filters that enhance plate features while suppressing distractors. Filtered features are then processed through a 2D Mamba block [18] for context modeling. This design improves robustness under low contrast, occlusion, or interference, directly addressing the challenges of weak feature perception and fine-grained structural semantics. Third, we embed the density-aware reordering state space module (DRSSM) at the tail of MDCNeXt to refine predictions in densely packed regions. General sequential scanning in SSM often mixes background and foreground features, fragmenting the already sparse electrode signals. This problem worsens in samples with densely packed plates, where spatial entanglement further weakens class-specific representations. DRSSM addresses this by using the coarse point map to semantically regroup features (e.g., cathode, anode, background) and reorder them into contiguous sequences. These reordered tokens are processed via state space modeling to enhance intra-class consistency. The refined features are then mapped back to their original layout, yielding sharper boundary localization and better category separation. Last, we design an adaptive label generation strategy for point segmentation. Specifically, we compute the spatial distance between adjacent homopolar plates based on annotations and use it as a reference radius to generate scale-aware ground truth masks. Unlike fixed-size masks, this adaptive mask better matches real spacing variations, improving robustness in both sparse and dense plate regions.
@@
  • We compare MDCNeXt with the corner detection, crowd counting, general/tiny object detection and image segmentation methods. Extensive experiments demonstrate that the proposed MDCNeXt performs favorably against the state-of-the-art methods under different metrics.
@@
This paper is based on and extends our CVPR version [21] in the following aspects. I) We significantly expand the dataset from 1,500 to 5,000 images, forming a new benchmark PBD5K with richer variations in image resolution, plate number, overhang distribution and interference types, and provide a more detailed dataset analysis. II) An intelligent annotation pipeline is introduced to improve labeling efficiency and consistency, providing the industrial AI community with a transparent and reproducible workflow. III) We upgrade the MDCNet [21] to a stronger version named MDCNeXt by embedding the proposed prompt-filtered state space module and density-aware reordering state space module. IV) We conduct more extensive experiments, including quantitatively and qualitatively comparisons with generalist, general and specialized segmentation models in terms of ten metrics, showing consistent superiority of MDCNeXt. V) We provide a comprehensive introduction to related work, detailed implementation details, and thorough ablation studies. VI) We present an in-depth discussion on future directions of PBD, outlining promising research opportunities.
II Related Work
II-A Open Vision Problem Modeling
@@
Emerging open vision problems often deviate from traditional tasks due to novel instance definitions, annotation challenges, or perceptual ambiguity, demanding tailored task abstraction and modeling. Liu et al. [4] formulate Chinese character stroke extraction as stroke-level instance segmentation, enabling fine-grained and transferable structural decomposition. Zhou et al. [5, 22] address the challenge of insubstantial object detection by treating faint or transparent targets as spatially diffuse regions with geometric uncertainty, and model detection through 3D voxel shift field estimation. Pang et al. [23] define open-vocabulary camouflaged object segmentation as a hybrid of open-set recognition and camouflage-aware segmentation, integrating vision-language priors and structural clues like depth and edges to capture imperceptible objects. Chiu et al. [24] frame wire segmentation as a thin-structure semantic segmentation task under ultra-high-resolution constraints, emphasizing continuity-preserving modeling tailored for elongated, slender targets. Yang et al. [25] conceptualize traffic sign interpretation as a semantic reasoning task over interrelated signs, where the goal is to interpret visual signals into natural-language instructions through multi-level relational logic. Zhou et al. [26] tackle the nighttime optical flow estimation problem by casting it as cross-domain motion alignment, leveraging reflectance and gradient-aligned latent spaces to transfer motion priors from auxiliary domains under degraded visibility. Liu et al. [27] design technical drawing parsing as a vector-based structural graph learning problem, transforming CAD designs into spatial graphs where geometric primitives are modeled as relational nodes for precise topological inference. Wei et al. [28] further leverage a type-agnostic and expressive line-based representation of graphical primitives, instead of traditional point-based methods, enabling a unified abstraction that accommodates diverse elements in engineering and architectural diagrams.
@@
II-B Object Detection, Counting and Segmentation
@@
Corner detection is widely used in motion detection, image matching, and video tracking. A corner point is typically defined as the intersection of two edges and can also be referred to as a feature point. Harris and Stephens [29] propose one of the earliest methods by directly computing the differential of the corner score with respect to direction. Shi–Tomasi [30] improves stability by selecting corners based on the minimum eigenvalue of the gradient matrix. Sub-pixel methods [31] further enhance accuracy by refining corner positions to real coordinates for geometric measurement or calibration. However, corner detectors are often not robust and typically require high redundancy to mitigate the impact of individual errors on recognition tasks.
@@
Object detection aims to locate and classify semantic objects in images or videos. Many methods rely on bounding box prediction, using rectangles to approximate object locations. Detectors are typically categorized into two paradigms: I) Two-stage detectors, such as Fast RCNN [32], Faster R-CNN [33], FPN [10], and RFCN [34], which generate region proposals before classification and refinement. II) One-stage detectors, such as RetinaNet [35], YOLOv10 [36], FCOS [37], and DETR [38], which directly predict bounding boxes and categories in a single forward pass. Although most single-stage methods have fast detection speed, they are usually not as accurate as the anchor-based two-stage detectors, especially in small objects. To bridge this gap, a number of tiny object detection methods have been developed to achieve a trade-off between accuracy and efficiency, including C3Det [39], CFINet [40], and HS-FPN [41].
@@
Dense object counting task is to infer the number of objects in dense scenarios. Most CNNs-based methods [42, 43, 44, 45] predict a density map from a crowd image, where the summation of the density map is the crowd count. For the structures, some works are proposed to address scale variation [44, 46, 47], to refine the predicted density map [45, 48, 49], and to encode context information [50, 51]. For the loss function, L2 loss is usually used in many models but it is sensitive to the choice of variance in the Gaussian kernel [52]. Therefore, Bayesian loss [53] is proposed with point-wise supervision.
@@
Image segmentation aims to simplify or transform an image into a more structured representation, facilitating further analysis. Diverse segmentation types exist in practice, including salient, camouflaged, transparent, and medical segmentation. Most methods [54, 55, 56] adopt a U-shaped encoder-decoder architecture [11, 10], where multi-level features are progressively fused to generate high-resolution masks. To enhance feature representation, some works [57, 58, 59, 60] incorporate attention mechanisms, such as spatial, channel-wise and gated attention to selectively emphasize informative clues. Meanwhile, handling scale variation among objects remains a key challenge. This has motivated research into multi-scale feature extraction techniques [61, 62, 63, 64, 65], including ASPP [62], Fold-ASPP [59], and DenseASPP [63], which aim to capture semantic clues across receptive fields of varying sizes. Recently, Transformer-based models [66, 67, 68] have attracted attention for their ability to model global context. Unlike CNNs with limited receptive fields, Transformers [69] utilize self-attention to capture long-range dependencies, improving performance in complex spatial tasks. Segformer [67] and SwinUNet [68] couple hierarchical encoding with attention-based decoding. However, their quadratic complexity and weak local bias limit scalability and fine-grained prediction. In contrast, Mamba-based methods [70, 71] leverage selective state-space models to replace attention with structured recurrence, enabling linear time complexity and superior scalability. Mamba [17] excels in global dependency modeling with low memory cost and an inherent structural bias, suiting thin-structure or topology-aware tasks. Some hybrid CNN-Mamba models [72, 73] combine local detail preservation with global reasoning, and providing an efficient framework for high-resolution and real-time segmentation. With the growing pursuit of foundation models and artificial general intelligence, unified and generalist segmentation models are emerging to address diverse segmentation tasks with a single set of parameters. SAM [74] is trained on 1.1 billion masks, enabling high-quality segmentation with various prompts. UniverSeg [75] focuses on unifying medical image segmentation across diverse organs and modalities. SegGPT [76] offers flexible segmentation via visual in-context learning. HQSAM [77] improves SAM for high-resolution segmentation. SAM 2 [78] extends SAM to video by incorporating memory attention and multi-frame prompts. For complex context-dependent concepts, EVP [79] and GateNetv2 [61] integrate low-level structural prompts and gated context, while Spider [80] and VSCode [81] apply 2D prompt learning to model foreground-background relations. In this work, our proposed MDCNeXt integrates the aforementioned advanced techniques, including CNN-Mamba hybrid design, attention-based refinement, multi-scale modeling, and prompt-driven filtering, aiming to provide a high-precision segmentation for the PBD task.
II-C Enhancing Weak Features in Complex Vision Tasks

### v2

@@ Abstract
Power batteries are essential components in electric vehicles, where internal structural defects can pose serious safety risks. We conduct a comprehensive study on a new task, power battery detection (PBD), which aims to localize the dense endpoints of cathode and anode plates from industrial X-ray images for quality inspection. Manual inspection is inefficient and error-prone, while traditional vision algorithms struggle with densely packed plates, low contrast, scale variation, and imaging artifacts. To address this issue and drive more attention into this meaningful task, we present PBD5K, the first large-scale benchmark for this task, consisting of 5,000 X-ray images from nine battery types with fine-grained annotations and eight types of real-world visual interference. To support scalable and consistent labeling, we develop an intelligent annotation pipeline that combines image filtering, model-assisted pre-labeling, cross-verification, and layered quality evaluation. We formulate PBD as a point-level segmentation problem and propose MDCNeXt, a model designed to extract and integrate multi-dimensional structure clues including point, line, and count information from the plate itself. To improve discrimination between plates and suppress visual interference, MDCNeXt incorporates two state space modules. The first is a prompt-filtered module that learns contrastive relationships guided by task-specific prompts. The second is a density-aware reordering module that refines segmentation in regions with high plate density. In addition, we propose a distance-adaptive mask generation strategy to provide robust supervision under varying spatial distributions of anode and cathode positions. Without any bells and whistles, our segmentation-based MDCNeXt consistently outperforms various other corner detection, crowd counting and general/tiny object detection-based solutions, making it a strong baseline that can help facilitate future research in PBD. Finally, we provide a detailed discussion of future directions and challenges for research in power battery detection.
Figure 1: Motivation of power battery detection (PBD). (a) EV sales are rapidly increasing, raising battery safety demands. (b) Battery packs account for $\sim$35% of EV cost, highlighting their critical role. (c) Assembly process of power batteries for EVs, where PBD is applied to the battery cells before assembly. (d) Illustration of the power battery detection task.
1 Introduction
@@
With the rapid adoption of electric vehicles (EVs), Fig. 1(a) shows that EV market share has grown from around 1% to over 20% in the past five years, with 17.1 million units sold globally in 2024 evsales2024 ; ieaev2024 . The power battery serves as the sole energy source for EVs, directly influencing their performance, range, and safety BEV_1 . As shown in Fig. 1(b), the battery pack contributes nearly 35% of total manufacturing cost, underscoring its critical economic and functional role. Ensuring its structural integrity requires reliable inspection throughout production. During cell assembly (Fig. 1(c)), hundreds of cathode and anode plates are stacked alternately. Any misalignment or overhang can lead to short circuits, overheating, or even explosion, creating serious safety risks. To address this, a post-assembly inspection process, termed power battery detection (PBD), is widely applied. As shown in Fig. 1(d), PBD uses X-ray images captured by digital radiography (DR) systems to localize cathode and anode endpoints with coordinate-level precision. This enables automatic computation of key quality indicators, such as the number of plates and their overhang [^1] Number and overhang are two critical metrics in battery quality control. Number indicates the total count of anode/cathode plates, while overhang measures the average vertical offset between one anode and its adjacent cathodes., which determine whether a battery cell is acceptable (OK) or defective (No Good, NG). However, long-term manual inspections are subject to visual fatigue and human errors. Even top manufacturers relying on large re-inspection teams face high labor costs and limited efficiency.
@@
Figure 2: Visual examples from the PBD5K dataset (best viewed zoomed in). (a) 1-pixel localization on high-resolution (3000 × 4000) X-ray images. (b) The imaging field-of-view include Close Shot (CS), Medium Shot (MS) and Long Shot (LS). (c) Examples of various attributes. See Tab. 1 for details.
Table 1: Attribute descriptions (see examples in Fig. 2(c)).
@@
Therefore, the development of an intelligent PBD model is an imminent need. To benchmark this new task, we first introduce PBD5K, the first large-scale power battery dataset with 5,000 X-ray images. It is constructed using an intelligent annotation pipeline that combines automatic filtering, model-assisted labeling, cross-validation, and hierarchical quality control, significantly improving labeling efficiency and accuracy. Compared with traditional visual tasks, PBD presents several unique challenges: I) Open Vision Problem Modeling. Like emerging tasks such as character stroke extraction CCSE , insubstantial object detection IOD , open vocabulary object detection and segmentation OVD ; OVSeg , how to model a suitable AI-based solution is critical for the PBD task. II) Single Pixel-Level Object Localization. As shown in Fig. 2(a), PBD requires accurately identifying plate endpoints that are only one pixel in size among millions of pixels, and providing their precise coordinates. This level of search difficulty is akin to finding a needle in a haystack. III) Multi-shot Imaging. As shown in Fig. 2(b), X-ray imaging is conducted under various field-of-view settings, leading to significant variations in spatial resolution. Even under the same imaging setup, batteries may exhibit different physical dimensions or placement scales. IV) Weak Feature Perception. As clearly observed in Fig. 2(c), the plate endpoints exhibit extremely weak features and are visually similar to their surrounding regions. Especially under interference from bifurcated plates, separators, or trays, their visibility is further diminished. V) Fine-grained Structural Semantics. PBD5K contains a wide range of overhang distribution patterns. When the distance between adjacent cathode and anode plates is small (less than 100 pixels), their visual appearances become easily confused. Since power batteries are manufactured with strict alternation between cathode and anode plates, accurate endpoint localization implicitly requires the model to understand the structural semantics, including the classification and ordering of electrode plates.
@@
In this paper, we propose a multi-dimensional collaborative network for accurate PBD. Imitating the coarse-to-fine human visual perception system coarse-to-fine1 ; coarse-to-fine2 , the model first perceives the battery region and then gradually refines the localization of each plate. Once the endpoints are precisely segmented, both their coordinates and total numbers can be derived. First, we reformulate PBD as a point-centric segmentation task, where the core objective is to predict a point map of single-pixel endpoints for cathode and anode plates. A U-shaped encoder-decoder FPN ; Unet serves as the basic structure for point prediction. Inspired by multi-task learning Hard_parameter_sharing ; soft_parameter_sharing ; MMFT and the structural traits of power batteries, we introduce two auxiliary tasks: line segmentation and plate counting. As shown in Fig. 3, connecting predicted endpoints yields line maps that constrain spatial position and reduce ambiguity. Line supervision guides endpoint refinement, while the counting branch enforces consistency between the predicted and actual number of electrodes, mitigating false positives and omissions. By jointly learning from these three complementary clues, we establish a segmentation-centered multi-dimensional collaborative framework. Next, to improve localization under weak textures and noisy backgrounds, we utilize state space models (SSMs) LSSL ; S4 ; Mamba ; VMamba , which offer efficient long-range modeling with linear complexity. Compared to Transformer-based models ViT ; Swin , SSMs are more suitable for high-resolution industrial images due to their lower computational cost. However, direct use of SSMs may amplify irrelevant large-region patterns (e.g., trays or tabs), thereby overwhelming the subtle and sparse endpoint signals. To address this, we propose the prompt-filtered state space module (PFSSM). It uses features from a representative prompt image to generate dynamic filters that enhance plate features while suppressing distractors. Filtered features are then processed through a 2D Mamba block VMamba for context modeling. This design improves robustness under low contrast, occlusion, or interference, directly addressing the challenges of weak feature perception and fine-grained structural semantics. Third, we embed the density-aware reordering state space module (DRSSM) at the tail of MDCNeXt to refine predictions in densely packed regions. General sequential scanning in SSM often mixes background and foreground features, fragmenting the already sparse electrode signals. This problem worsens in samples with densely packed plates, where spatial entanglement further weakens class-specific representations. DRSSM addresses this by using the coarse point map to semantically regroup features (e.g., cathode, anode, background) and reorder them into contiguous sequences. These reordered tokens are processed via state space modeling to enhance intra-class consistency. The refined features are then mapped back to their original layout, yielding sharper boundary localization and better category separation. Last, we design an adaptive label generation strategy for point segmentation. Specifically, we compute the spatial distance between adjacent homopolar plates based on annotations and use it as a reference radius to generate scale-aware ground truth masks. Unlike fixed-size masks, this adaptive mask better matches real spacing variations, improving robustness in both sparse and dense plate regions.
@@
  • We compare MDCNeXt with corner detection, crowd counting, general/tiny object detection and image segmentation methods. Extensive experiments demonstrate that the proposed MDCNeXt performs favorably against the state-of-the-art methods under different metrics.
@@
This paper is based on and extends our CVPR version MDCNet in the following aspects. I) We significantly expand the dataset from 1,500 to 5,000 images, forming a new benchmark PBD5K with richer variations in image resolution, plate number, overhang distribution and interference types, and provide a more detailed dataset analysis. II) An intelligent annotation pipeline is introduced to improve labeling efficiency and consistency, providing the industrial AI community with a transparent and reproducible workflow. III) We upgrade the MDCNet MDCNet to a stronger version named MDCNeXt by embedding the proposed prompt-filtered state space module and density-aware reordering state space module. IV) We conduct more extensive experiments, including quantitatively and qualitatively comparisons with generalist, general and specialized segmentation models in terms of ten metrics, showing consistent superiority of MDCNeXt. V) We provide a comprehensive introduction to related work, detailed implementation details, and thorough ablation studies. VI) We present an in-depth discussion on future directions of PBD, outlining promising research opportunities.
2 Related Work
2.1 Open Vision Problem Modeling
@@
Emerging open vision problems often deviate from traditional tasks due to novel instance definitions, annotation challenges, or perceptual ambiguity, demanding tailored task abstraction and modeling. Liu et al. CCSE formulate Chinese character stroke extraction as stroke-level instance segmentation, enabling fine-grained and transferable structural decomposition. Zhou et al. IOD ; GOD address the challenge of insubstantial object detection by treating faint or transparent targets as spatially diffuse regions with geometric uncertainty, and model detection through 3D voxel shift field estimation. Pang et al. OVCOS define open-vocabulary camouflaged object segmentation as a hybrid of open-set recognition and camouflage-aware segmentation, integrating vision-language priors and structural clues like depth and edges to capture imperceptible objects. Chiu et al. WireSegHR frame wire segmentation as a thin-structure semantic segmentation task under ultra-high-resolution constraints, emphasizing continuity-preserving modeling tailored for elongated, slender targets. Yang et al. TSI conceptualize traffic sign interpretation as a semantic reasoning task over interrelated signs, where the goal is to interpret visual signals into natural-language instructions through multi-level relational logic. Zhou et al. NighttimeOpticalFlow tackle the nighttime optical flow estimation problem by casting it as cross-domain motion alignment, leveraging reflectance and gradient-aligned latent spaces to transfer motion priors from auxiliary domains under degraded visibility. Liu et al. VectorGraphNET design technical drawing parsing as a vector-based structural graph learning problem, transforming CAD designs into spatial graphs where geometric primitives are modeled as relational nodes for precise topological inference. Wei et al. VecFormer further leverage a type-agnostic and expressive line-based representation of graphical primitives, instead of traditional point-based methods, enabling a unified abstraction that accommodates diverse elements in engineering and architectural diagrams.
@@
2.2 Object Detection, Counting and Segmentation
@@
Corner detection is widely used in motion detection, image matching, and video tracking. A corner point is typically defined as the intersection of two edges and can also be referred to as a feature point. Harris and Stephens Harris propose one of the earliest methods by directly computing the differential of the corner score with respect to direction. Shi–Tomasi Shitomasi improves stability by selecting corners based on the minimum eigenvalue of the gradient matrix. Sub-pixel methods Sub-Pixel further enhance accuracy by refining corner positions to real coordinates for geometric measurement or calibration. However, corner detectors are often not robust and typically require high redundancy to mitigate the impact of individual errors on recognition tasks.
@@
Object detection aims to locate and classify semantic objects in images or videos. Many methods rely on bounding box prediction, using rectangles to approximate object locations. Detectors are typically categorized into two paradigms: I) Two-stage detectors, such as Fast RCNN Fast_rcnn , Faster R-CNN Faster_rcnn , FPN FPN , and RFCN RFCN , which generate region proposals before classification and refinement. II) One-stage detectors, such as RetinaNet RetinaNet , YOLOv10 Yolov10 , FCOS FCOS , and DETR DETR , which directly predict bounding boxes and categories in a single forward pass. Although most single-stage methods have fast detection speed, they are usually not as accurate as the anchor-based two-stage detectors, especially in small objects. To bridge this gap, a number of tiny object detection methods have been developed to achieve a trade-off between accuracy and efficiency, including C3Det C3Det , CFINet CFINet , and HS-FPN HS-FPN .
@@
Dense object counting task is to infer the number of objects in dense scenarios. Most CNNs-based methods DM ; CUT_cc ; scale_variation_cc1 ; density_map_cc1 predict a density map from a crowd image, where the summation of the density map is the crowd count. For the structures, some works are proposed to address scale variation scale_variation_cc1 ; scale_variation_cc2 ; scale_variation_cc3 , to refine the predicted density map density_map_cc1 ; density_map_cc2 ; IOCFormer , and to encode context information context_information_cc1 ; context_information_cc2 . For the loss function, L2 loss is usually used in many models but it is sensitive to the choice of variance in the Gaussian kernel Adaptive_CC . Therefore, Bayesian loss BL is proposed with point-wise supervision.
@@
Image segmentation aims to simplify or transform an image into a more structured representation, facilitating further analysis. Diverse segmentation types exist in practice, including salient, camouflaged, transparent, and medical segmentation. Most methods DeepLabV3+ ; MINet ; MSNet_Polyp adopt a U-shaped encoder-decoder architecture Unet ; FPN , where multi-level features are progressively fused to generate high-resolution masks. To enhance feature representation, some works AMP ; BDRAR_Shadow ; GateNet ; li2023delving incorporate attention mechanisms, such as spatial, channel-wise and gated attention to selectively emphasize informative clues. Meanwhile, handling scale variation among objects remains a key challenge. This has motivated research into multi-scale feature extraction techniques GateNetv2 ; ASPP ; DenseASPP ; M2SNet ; ZoomNet , including ASPP ASPP , Fold-ASPP GateNet , and DenseASPP DenseASPP , which aim to capture semantic clues across receptive fields of varying sizes. Recently, Transformer-based models Segmenter ; SegFormer ; SwinUNet have attracted attention for their ability to model global context. Unlike CNNs with limited receptive fields, Transformers transformer utilize self-attention to capture long-range dependencies, improving performance in complex spatial tasks. Segformer SegFormer and SwinUNet SwinUNet couple hierarchical encoding with attention-based decoding. However, their quadratic complexity and weak local bias limit scalability and fine-grained prediction. In contrast, Mamba-based methods Sigma ; Segmamba leverage selective state-space models to replace attention with structured recurrence, enabling linear time complexity and superior scalability. Mamba Mamba excels in global dependency modeling with low memory cost and an inherent structural bias, suiting thin-structure or topology-aware tasks. Some hybrid CNN-Mamba models U-mamba ; CM-UNet combine local detail preservation with global reasoning, and providing an efficient framework for high-resolution and real-time segmentation. With the growing pursuit of foundation models and artificial general intelligence, unified and generalist segmentation models are emerging to address diverse segmentation tasks with a single set of parameters. SAM SAM is trained on 1.1 billion masks, enabling high-quality segmentation with various prompts. UniverSeg UniverSeg focuses on unifying medical image segmentation across diverse organs and modalities. SegGPT SegGPT offers flexible segmentation via visual in-context learning. HQSAM HQSAM improves SAM for high-resolution segmentation. SAM 2 SAM2 extends SAM to video by incorporating memory attention and multi-frame prompts. For complex context-dependent concepts, EVP EVP and GateNetv2 GateNetv2 integrate low-level structural prompts and gated context, while Spider CDCU-Spider and VSCode VSCode apply 2D prompt learning to model foreground-background relations. In this work, our proposed MDCNeXt integrates the aforementioned advanced techniques, including CNN-Mamba hybrid design, attention-based refinement, multi-scale modeling, and prompt-driven filtering, aiming to provide a high-precision segmentation for the PBD task.
2.3 Enhancing Weak Features in Complex Vision Tasks